OpenStudy (anonymous):

Find the absolute maximum and absolute minimum values of f on the given interval. f(t) = 2 cos t + sin 2t, [0, π/2]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

-2sint+2cos2t is the derivativ>

ganeshie8 (ganeshie8):

yes, next what

OpenStudy (anonymous):

then i convern the 2(cos2t) into sin....so would that be 1-sin^2(sx)

OpenStudy (anonymous):

sin^2(2x) i mean

ganeshie8 (ganeshie8):

set the first derivative equal to 0 and solve \(t\) in the given interval

OpenStudy (anonymous):

but how

ganeshie8 (ganeshie8):

\(\large -2\sin t+2\cos(2t) = 0\)

ganeshie8 (ganeshie8):

we use this trig identity : \(\large \cos (2x) = 1-2\sin^2x\)

OpenStudy (anonymous):

-sint+cos(2t)=0

OpenStudy (anonymous):

oh ok

ganeshie8 (ganeshie8):

\(\large -2\sin t+2\cos(2t) = 0\) \(\large -\sin t+\cos(2t) = 0\) \(\large -\sin t+1-2\sin^2t = 0\) \(\large 2\sin^2t + \sin t -1= 0\)

ganeshie8 (ganeshie8):

see if u can factor it

OpenStudy (anonymous):

sin t = -1 sin t =1 ?

OpenStudy (anonymous):

so - pi/2 and positive pi/2 but only positive because its in range?

ganeshie8 (ganeshie8):

\(\large 2\sin^2t + \sin t -1= 0\) \(\large 2x^2 + x - 1= 0\) \(\large 2x^2 + 2x-x - 1= 0\) \(\large 2x(x + 1)-1(x + 1)= 0\) \(\large (x+1)(2x-1)= 0\) x = -1, 1/2 \(\large \sin t = -1\) or \(\large \sin t = 1/2\)

ganeshie8 (ganeshie8):

\(\large t = \pi/6\)

OpenStudy (anonymous):

oh ok i see

ganeshie8 (ganeshie8):

So you need to test your function at \(t = \pi/6\) and the boundary values

ganeshie8 (ganeshie8):

\(\large f(t) = 2 \cos t + \sin (2t)\) f(pi/6) = ? f(0) = ? f(pi/2) = ?

ganeshie8 (ganeshie8):

maximum of above 3 values is the `absolute maximum` in that interval minimum of above 3 values is the `absolute minimum` in that interval

OpenStudy (anonymous):

thank you

ganeshie8 (ganeshie8):

yw :)